# Generalized Spectral Clustering of Low-Inertia Power Networks

### v1

@@ Abstract
Large-scale integration of distributed energy resources has led to a rapid increase in the number of controllable devices and a significant change in system dynamics. This has necessitating the shift towards more distributed and scalable control strategies to manage the increasing system complexity. In this work, we address the problem of partitioning a low-inertia power network into dynamically coherent subsystems to facilitate the utilization of distributed control schemes. We show that an embedding of the power network using the spectrum of the linearized synchronization dynamics matrix results in a natural decomposition of the network. We establish the connection between our approach and the broader framework of spectral clustering using the Laplacian matrix of the admittance network. The proposed method is demonstrated on the IEEE 30-bus test system, and numerical simulations show that the resulting clusters using our approach are dynamically coherent. We consider the robustness of the clusters identified in the network by analyzing the sensitivity of the small eigenvalues and their corresponding eigenspaces, which determines the coherency structure of the oscillator dynamics, to variations in the steady-state operating points of the network.
@@ I Introduction
The transition to a more sustainable energy system typically requires the integration of large quantities of renewable generation sources, micro-generators, storage, and responsive loads engaged in the real-time demand and supply balancing at different levels of the grid. A significant portion of these distributed energy resources (DERs) are small-scale – less than $1$ MW. With recent revision of regulatory measures, such as FERC order $2222$ [21], enabling the direct participation of DERs in electricity markets, the grid of the future will be composed of a large number of these controllable resources that require real-time coordination. In addition, the gradual replacement of synchronous generators – which are dispatchable and contribute to the system inertia – with intermittent and variable generation from renewables, leads to a gradual loss of the robust frequency and voltage control inherent to synchronous machines[16]. The need to manage this increasingly complex system has necessitated the development of different operational paradigms, including optimization-based control schemes for automatic generation control [41], which have been shown to be fast-acting and spatially precise, enabling coordinated response to system disturbances.
@@
In most modern power systems, the independent system operator (ISO) is responsible for the centralized coordination of all system resources. However, as the number of decision variables and operational constraints increases, solving the real-time centralized coordination and control problem can become computationally prohibitive. Furthermore, the communication requirement between a centralized controller and these resources poses an additional challenge to scalability.
@@
Several works [45, 24, 8] have shown the potential of distributed control schemes – such as distributed model-predictive control[41, 42], multi-agent systems[27, 9], and consensus-based methods[7, 15] – for achieving acceptable trade-offs between algorithmic performance and communication/coordination requirements for solving centralized control problems. Chanfreut et al. [8] further showed that this balance between the dual objectives of performance and communication requirements in distributed control schemes is analogous to that of system partitioning, which seeks to identify a suitable decomposition of a global system into suitable subsystems. In the context of networked control systems, this decomposition can be framed as a network partitioning problem, where the partitions should be such that the control actions and disturbances originating within each subsystem are reasonably confined to that partition and have minimal impact on the rest of the network. In this work, the term partition is used synonymously with cluster, control zone, subsystem, and control area.
@@
In conventional power networks (like PJM), the current zones and zonal boundaries in the network result from historical asset ownership rather than the properties of the network [10]. Control zones defined in this way may be sufficient for system planning and market operations but are not necessarily suited for distributed and decentralized power system operations. To address this need, recent works have approached the problem of network partitioning in power networks with different objectives including contingency management [37, 13, 31], generator coherency identification [32, 46], designing localized reactive power markets [47], localized voltage control [25, 29], and parallel power system restoration [22] using methods including spectral clustering techniques [30, 34, 13, 44], evolutionary computation algorithms [12, 6], information theoretic approaches [25, 5], and sensitivity analysis techniques [38, 29].
In this work, we address the problem of partitioning power networks into control areas to facilitate scalable distributed control strategies. We extend the spectral clustering approach developed in [34, 30], where the spectrum of a graph whose edges represent branch susceptances is used to identify clusters in a power network. Our approach uses the spectrum of the linearized non-uniform Kuramoto oscillator dynamics matrix to construct a low-dimensional embedding of the network. We show that the spectrum of this matrix is the same as the spectrum of a generalized eigenvalue problem defined on a graph whose edge weights represent the sensitivity of the real power flow on the lines to line angles, with node weights representing the effective damping of each oscillator. This framework allows us to incorporate the effect of the steady-state voltage magnitudes, angles, and node types (reflected in time constants of the oscillators) in the clustering problem.
The proposed approach is demonstrated on the IEEE 30-bus network for various numbers of clusters. Numerical simulations show that the identified clusters are dynamically coherent with perturbations originating within each cluster largely confined to that cluster. These coherent groups of oscillators, identified using the linearized oscillator dynamics equation, are shown to be robust to changes in the network operating condition around the nominal state, provided the relative spectral gap is large. Additionally, we provide bounds on the number of clusters and the stability of the resulting network embedding.
@@
The remainder of the paper is organized as follows: Section II establishes some graph theory and power system preliminaries, Section III describes the clustering optimization problem and the generalized spectral clustering solution. Section IV presents results on a test power system, with simulations of the impact of disturbances on the network dynamics and robustness analysis of the resulting clusters presented in Section V. Finally, Section VI concludes the paper.
@@ II-A Graph Theory Preliminaries
We define an undirected graph $\mathcal{G}$ as the triple $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$, where $\mathcal{V}=\{1,\ldots,n\}$ is the set of vertices (nodes), $|\mathcal{V}|=n$ is the number of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges (branches), $|\mathcal{E}|=m$ is the number of edges, and $W\in\mathbb{R}^{n\times n}$ is the weighted symmetric adjacency matrix of the graph. $\{i,j\}\in\mathcal{E}$ represents the undirected edge connecting nodes $i$ and $j$, and $w_{ij}\in\mathbb{R}_{>0}$ is the corresponding weight of the edge. The graph $\mathcal{G}$ is simple, if $w_{ii}=0$ for all $i$ (i.e., there are no self-loops) and a single edge connects every pair of connected nodes. The graph $\mathcal{G}$ is connected if there is a path between every pair of distinct nodes in $\mathcal{V}$.
@@
If we assign a unique number $\ell\in\{1,\ldots,m\}$ and an arbitrary direction to each edge $\{i,j\}\in\mathcal{E}$, we define the oriented incidence matrix $B\in\mathbb{R}^{n\times m}$ element-wise as $B_{k\ell}=1$ if node $k$ is a sink node of the edge $\ell$, 1-1 if node $k$ is a source node of edge $\ell$, and 0 otherwise. Then for a vector $x\in\mathbb{R}^{n}$, the vector $B^{\intercal}x$ has entries $x_{i}-x_{j}$ corresponding to each branch $\ell\in\mathcal{E}$ connecting node $i$ and $j$.
The Laplacian matrix of the graph $\mathcal{G}$ denoted by $L\in\mathbb{R}^{n\times n}$ is the symmetric positive semi-definite (PSD) matrix defined as L=DWL=D-W, where $D=\text{diag}\left([d_{1},\ldots,d_{n}]^{\intercal}\right)$ is the diagonal degree matrix and $d_{i}=\sum_{j}w_{ij}$ is the degree of node $i$. The $ij$-th element of $L$
$L_{ij}=\begin{cases}d_{i}&\text{if }i=j,\\ -w_{ij}&\text{if }\{i,j\}\in\mathcal{E},\\ 0&\text{otherwise}.\end{cases}$
@@
We denote the positive definiteness and semi-definiteness of a symmetric matrix $M=M^{\intercal}$ as $M\succ 0$ and $M\succeq 0$, respectively. $\mathbf{1}_{n}\in\mathbb{R}^{n}$ and $\mathbf{0}_{n}\in\mathbb{R}^{n}$ denotes the $n$-dimensional vectors of ones and zeros, respectively, $\text{Ker}(L)=\text{Ker}(B^{\intercal})=\text{Span}(\mathbf{1}_{n})$, that is $L\mathbf{1}_{n}=\mathbf{0}_{n}$, this subspace is commonly referred to as the agreement subspace of $\mathcal{G}$ [1]. Let $\lambda_{i}$ denote the $i$th smallest eigenvalue of the matrix $L$, since the Laplacian is symmetric and positive semi-definite, its eigenvalues can be ordered as as $0=\lambda_{1}\leq\lambda_{2}\leq\ldots\leq\lambda_{n}$. If the graph $\mathcal{G}$ is a connected graph, the second smallest eigenvalue, known as the algebraic connectivity $\lambda_{2}>0$.
@@
Given a pair of matrices $L,D\in\mathbb{R}^{n\times n}$, the nonzero vector $v\in\mathbb{R}^{n}$ is called an eigenvector of the pair $(L,D)$ if there are scalars $\alpha,\beta\in\mathbb{R}$, not both zero such that $\alpha Lv=\beta Dv$. The scalar $\lambda(L,D)=\alpha/\beta$ is the eigenvalue associated with the eigenvector $v$. We denote by $\lambda_{i}(L,D)$ the $i$th smallest eigenvalue of the matrix pair $(L,D)$.
For a non-empty set $\mathcal{V}_{1}\subset\mathcal{V}$, we denote the subgraph induced by $\mathcal{V}_{1}$ on $\mathcal{G}$ as $\mathcal{G}_{\mathcal{V}_{1}}=(\mathcal{V}_{1},\mathcal{E}_{1},W_{1})$ where $\mathcal{E}_{1}\subset\mathcal{E}\!\cap\!(\mathcal{V}_{1}\times\mathcal{V}_{1})$, and $W_{1}$ is the corresponding weighted adjacency matrix. We denote the boundary of the set $\mathcal{V}_{1}$ or the bound of the subgraph $\mathcal{G}_{\mathcal{V}_{1}}$ as
$\partial(\mathcal{V}_{1})=\sum_{i\in\mathcal{V}_{1},\,j\in\mathcal{V}_{1}^{c}}w_{ij}.$
@@ II-B Power System Preliminaries
In this work, we consider a lossless AC power network. Ignoring the effect of line charging (shunts) and assuming that the 3-phase system is balanced, the network can be modeled as an undirected weighted graph $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$, with $|\mathcal{V}|=n$ buses and $|\mathcal{E}|=m$ transmission lines. We assume that the graph is connected and that the set of nodes is given by $\mathcal{V}=\mathcal{V}_{G}\cup\mathcal{V}_{R}\cup\mathcal{V}_{L}$, where $\mathcal{V}_{G}$ is the set of buses with synchronous generators, $\mathcal{V}_{R}$ is the set of buses with renewable generation interfaced to the grid through inverters, and $\mathcal{V}_{L}$ is the set of load buses. The edge weight of the lossless line $\{i,j\}\in\mathcal{E}$ is given by $w_{ij}=\Im(Y_{ij})$, where $Y=Y^{\intercal}\in\mathbb{C}^{n\times n}$ is the complex admittance matrix (in per unit) of the network. Notice that the weighted admittance matrix $W=W^{\intercal}\in\mathbb{R}^{n\times n}$ for a lossless network.
We denote the per unit bus voltage magnitude at bus $i$ by $|V_{i}|\in\mathbb{R}$ and the voltage angle (in rad) as $\delta_{i}\in\mathbb{S}^{1}$, where $\mathbb{S}^{1}$ is the unit circle. Given $\delta\in\mathbb{T}^{n}=\mathbb{S}^{1}\times\mathbb{S}^{1}\times\cdots\times\mathbb{S}^{1}$ – the vector of bus voltage angles in the $n$-torus – the vector of line angles is defined as $\theta=B^{\intercal}\delta$, where $\theta\in\mathbb{T}^{m}$ and $B$ is the oriented incidence matrix of the network.
The real power flow on line $\{i,j\}\in\mathcal{E}$ can be expressed in terms of the bus voltage magnitudes, angles, and the entries of the admittance matrix as
$p_{ij}=-p_{ji}$ $=|V_{i}||V_{j}|\Im(Y_{ij})\sin(\delta_{i}-\delta_{j})$
$=|V_{i}||V_{j}|\Im(Y_{ij})\sin(\theta_{\ell})$
@@
where $M_{i}>0$ is the inertia coefficient of the synchronous generator connected at node $i\in\mathcal{V}_{G}$. The term $D_{i}>0$ represents the damping coefficient of the $i$th synchronous generator when $i\in\mathcal{V}_{G}$, the frequency-dependent component of the load connected to bus $i$ when $i\in\mathcal{V}_{L}$, and for droop-controlled inverter buses $i\in\mathcal{V}_{R}$, the total frequency dependence from the inverter’s droop and frequency-responsive loads (i.e., $R_{i}^{-1}+D_{i}^{\prime}$, where $R_{i}^{-1}$ is the droop coefficient and $D_{i}^{\prime}$ is the frequency-responsive component of the load) [1, 18]. $\omega_{i}$ is the natural frequency of the $i$th oscillator (this is the net real power injection at bus $i$) with $\omega_{i}>0$ for $i\in\mathcal{V}_{G}\cup\mathcal{V}_{R}$ and $\omega_{i}\leq 0$ for $i\in\mathcal{V}_{L}$.
With the integration of more renewable generation and the retirement of synchronous generation, we assume that the inertia coefficients $M_{i}$ are small compare damping term $D_{i}$ – this is typical for low inertia grids. [19] showed that under these assumptions, in the time-scale for synchronization, the network dynamics can be approximated by the dynamics of non-uniform Kuramoto oscillators where the model’s approximation error is $\mathcal{O}(\epsilon)$ which goes to zero as $t\rightarrow\infty$, $\epsilon$ is the singular perturbation parameter which typically is the worst-case choice of the ratio $M_{i}/D_{i}$.
The dynamics of the power network can then be reduced to the first-order non-uniform Kuramoto oscillator model
@@
When the connectivity of the network dominates the non-uniformity in the natural frequencies of the oscillators, a frequency-synchronized solution to (3) exists[17]. Let $|V^{*}|$ and $\delta^{*}$ denote the voltage magnitudes and angles of the frequency-synchronized network, the synchronous frequency of the oscillators denoted by $\dot{\delta}^{*}=\omega_{\text{sync}}\mathbf{1}_{n}$, where $\omega_{\text{sync}}=(\sum_{i}\omega_{i})/(\sum_{i}D_{i})$ and $\dot{\theta}^{*}=B^{\intercal}\dot{\delta}^{*}=\omega_{\text{sync}}(B^{\intercal}\mathbf{1}_{n})=\mathbf{0}_{m}$. In the synchronously rotating reference frame, the frequency-synchronized solution of (3) is an equilibrium of the system and the dynamic equations reduce to the AC power flow equations of a lossless power network given by
$\omega_{i}=\sum\limits_{j=1}^{n}|V_{i}^{*}||V_{j}^{*}|\Im(Y_{ij})\sin(\delta_{i}^{*}-\delta_{j}^{*}),\;i\in\mathcal{V}.$
II-C Linearized Dynamics
Assuming reactive power balance in the network can be controlled via local compensation, meaning that $|V_{i}^{*}|$ can be assumed to be constant (not necessarily $1.0$ p.u.), the linearized voltage angle dynamics around a locally stable synchronous solution $\delta^{*}$ is given by the weighted Laplacian flow, see appendix A for details,

### v2

@@ Abstract
Large-scale integration of distributed energy resources has led to a rapid increase in the number of controllable devices and a significant change in system dynamics. This has necessitating the shift towards more distributed and scalable control strategies to manage the increasing system complexity. In this work, we address the problem of partitioning a low-inertia power network into dynamically coherent subsystems to facilitate the utilization of distributed control schemes. We show that an embedding of the power network using the spectrum of the linearized synchronization dynamics matrix results in a natural decomposition of the network. We establish the connection between our approach and the broader framework of spectral clustering using the Laplacian matrix of the admittance network. The proposed method is demonstrated on the IEEE 30-bus test system. We consider the robustness of the clusters by analyzing the sensitivity of the small eigenvalues and their corresponding eigenspaces to perturbations caused by variation in the steady-state operating points of the network.
@@ I Introduction
The transition to a more sustainable energy system typically requires the large-scale integration of renewable generation resources, micro-generators, energy storage devices, and responsive loads, engaged in real-time balancing of supply and demand across different levels of the grid. Recent regulatory changes, such as FERC order $2222$ [19], enable direct participation of DERs in electricity markets, ensuring the grid of the future will be composed of a large number of these controllable resources that require real-time coordination. In addition, the gradual replacement of synchronous generators with intermittent and variable renewable generation leads to a gradual loss of the robust frequency and voltage control inherent to synchronous machines[15]. This increasingly complex system requires new operational paradigms, including optimization-based control schemes for automatic generation control [38], which have been shown to be fast-acting and spatially precise, enabling coordinated response to system disturbances.
@@
In most modern power systems, the independent system operator (ISO) is responsible for the centralized coordination of all system resources. However, as the number of decision variables and operational constraints increases, solving the real-time centralized coordination and control problem can become computationally prohibitive. Furthermore, the communication requirement between the centralized controller and these resources poses an additional scalability challenge.
@@
Several works [42, 22, 7] have shown the potential of distributed control schemes – such as distributed model-predictive control[38, 39], multi-agent systems[25, 8], and consensus-based methods[6, 14] for achieving acceptable trade-offs between algorithmic performance and communication/coordination requirements for centralized control problems. Chanfreut et al. [7] further showed that balancing the dual objectives of performance and communication requirements in distributed control schemes is analogous to that of system partitioning, which seeks to identify a suitable decomposition of a global system into suitable subsystems. In the context of networked control systems, this decomposition can be framed as a network partitioning problem, where the partitions should be such that the control actions and disturbances originating within each subsystem are reasonably confined to that partition, with minimal impact on the rest of the network. In this work, the term partition is used synonymously with cluster, control zone, subsystem, and control area.
@@
In conventional power networks (such as PJM), the current zones and zonal boundaries are determined by historical asset ownership rather than by the properties of the network [9]. Control zones defined in this way may be sufficient for system planning and markets, but are not necessarily suited for decentralized power system operations. To address this need, recent works have approached network partitioning in power networks with different objectives including contingency management [35, 12], generator coherency identification [30], designing localized reactive power markets [43], localized voltage control [23, 27], and parallel power system restoration [20] using methods including spectral clustering techniques [28, 32, 12, 41], evolutionary computation algorithms [11, 5], information theoretic approaches [23, 4], and sensitivity analysis techniques [36, 27].
In this work, we address the problem of partitioning power networks into control areas to facilitate scalable distributed control strategies. Some of the limitations of previous approaches [32, 37] are that they use a DC power flow model of the network, this approximation may fail to capture the true underlying network dynamics, especially when the phase angle differences become large. Additionally, many classic spectral clustering approaches for graphs [28] do not typically take into account node types, consequently, using them for clustering in power network can result in clusters without generator nodes. We extend the spectral clustering approach for power networks developed in [32], where the spectrum of the Laplacian of a graph whose edge weights represent branch susceptances is used to identify clusters in the network.
The main contributions of this paper are as follows:
  1. 1.
    We develop a generalized spectral clustering framework based on the linearized synchronization dynamics of low-inertia power networks, providing a dynamically meaningful decomposition of the system.
  2. 2.
    We formulate the clustering problem as a generalized eigenvalue problem that incorporates both network structure and heterogeneous damping, enabling the method to capture the influence of device-level dynamics on system behavior.
  3. 3.
    We provide conditions under which the identified clusters are robust to variations in the steady-state network operating condition.
  4. 4.
    We demonstrate this approach on the IEEE 30-bus network for various numbers of clusters, showing that the resulting clusters are dynamically coherent with perturbations largely confined to the originating cluster.
  5. 5.
    We provide a meaningful interpretation of the resulting clusters in the context of power systems.
@@
The remainder of the paper is organized as follows: Section II presents graph-theoretic and power-system preliminaries; Section III describes the clustering optimization problem and the generalized spectral clustering solution. Section IV presents results for a test power system, with simulations of the impact of disturbances on network dynamics and a robustness analyses of the resulting clusters presented in Section V. Finally, Section VI concludes the paper.
@@ II-A Graph Theory Preliminaries
Given a symmetric matrix $A$, the notation $A\succ 0$ and $A\succeq 0$ denotes positive definiteness and semi-definiteness, respectively. $\mathbf{1}_{n}\in\mathbb{R}^{n}$ and $\mathbf{0}_{n}\in\mathbb{R}^{n}$ denotes the $n$-dimensional vectors of all ones and zeros, respectively.
@@
We denote by $\mathcal{G}$ the undirected graph defined as the triple $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$, $\mathcal{V}=\{1,\ldots,n\}$ is the set of vertices and $|\mathcal{V}|=n$ is the number of vertices on the graph, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ denotes the edge set (i.e. branches) on the graph with $|\mathcal{E}|=m$ edges, and $W\in\mathbb{R}^{n\times n}$ denotes the weighted adjacency matrix. We use $\{i,j\}\in\mathcal{E}$ to represent the undirected edge connecting nodes $i,j\in\mathcal{V}$ with edge weight $w_{ij}\geq 0$. For a non-empty set $\mathcal{V}_{1}\subset\mathcal{V}$, we denote the boundary of the set $\mathcal{V}_{1}$ as $\partial(\mathcal{V}_{1})=\sum_{i\in\mathcal{V}_{1},\,j\in\overline{\mathcal{V}}_{1}}w_{ij}.$
@@
In this work, we assume the graph $\mathcal{G}$ is simple, that is, $w_{ii}=0$ for all $i$ and a single edge connects every pair of connected nodes. For each branch $\{i,j\}\in\mathcal{E}$, if we assign an arbitrary orientation and a unique number $\ell\in\{1,\ldots,m\}$, the oriented incidence matrix of the denoted by $B\in\mathbb{R}^{n\times m}$ is defined elementwise as $B_{k\ell}=1$ if node $k$ is a sink of the oriented edge $\ell$, 1-1 if node $k$ is a source for edge $\ell$, and 0 otherwise.
@@
The Laplacian matrix of the graph denoted by $L\in\mathbb{R}^{n\times n}$ is the symmetric positive semi-definite matrix defined as L=DWL=D-W, where $D$ is the diagonal matrix of node degrees $D_{ii}=\sum_{j}w_{ij}$ for each $i$. If we denote the $i$th smallest eigenvalue of $L$ by $\lambda_{i}$, the eigenvalues can be ordered as $0=\lambda_{1}\leq\lambda_{2}\leq\ldots\leq\lambda_{n}$. If the graph $\mathcal{G}$ is a connected, the second smallest eigenvalue (the algebraic connectivity) $\lambda_{2}>0$ and $\text{Ker}(L)=\text{Ker}(B^{\intercal})=\text{Span}(\mathbf{1}_{n})$, that is $L\mathbf{1}_{n}=\mathbf{0}_{n}$, this subspace is commonly referred to as the agreement subspace of $\mathcal{G}$ [1].
@@
Given a pair of symmetric matrices $L,D\in\mathbb{R}^{n\times n}$, we denote by $\lambda_{i}(L,D)$ the $i$th smallest generalized eigenvalue of the matrix pair $(L,D)$. $\lambda_{i}(L,D)\in\mathbb{R}$ is the scalar that solves the generalized eigenvalue equation $Lv=\lambda_{i}(L,D)Dv$, and the non-zero vector $v\in\mathbb{R}^{n}$ is the corresponding eigenvector.
@@ II-B Power System Preliminaries
In this work, we consider a lossless AC power system, that is, $r_{ij}=0$ for all $i$ and $j$. Ignoring the effect of line charging (shunts) and assuming that the 3-phase system is balanced, the network can be modeled as an undirected weighted graph $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$, with $|\mathcal{V}|=n$ nodes (buses) and $|\mathcal{E}|=m$ branches (transmission lines). We assume that the set of nodes can be partitioned as $\mathcal{V}=\mathcal{V}_{G}\cup\mathcal{V}_{R}\cup\mathcal{V}_{L}$, where $\mathcal{V}_{G}$ is the set of buses with synchronous generators, $\mathcal{V}_{R}$ is the set of buses with renewable generation connected to the grid via droop controlled inverters, and $\mathcal{V}_{L}$ is the set of load buses. The edge weight of the branch $\{i,j\}\in\mathcal{E}$ is given by $w_{ij}=\Im(Y_{ij})$, where $Y=Y^{\intercal}\in\mathbb{C}^{n\times n}$ is the complex admittance matrix (in per unit) of the network.
We denote the per unit bus voltage magnitude at bus $i$ by $|V_{i}|\in\mathbb{R}$ and the voltage angle (in rad) as $\delta_{i}\in\mathbb{S}^{1}$, where $\mathbb{S}^{1}$ is the unit circle. Given $\delta\in\mathbb{T}^{n}=\mathbb{S}^{1}\times\mathbb{S}^{1}\times\cdots\times\mathbb{S}^{1}$ – the vector of bus voltage phase angles in the $n$-torus – we define the vector of line angles as $\theta=B^{\intercal}\delta\in\mathbb{T}^{m}$, where $\theta_{\ell}=\delta_{i}-\delta_{j}$ for each edge $\ell\in\mathcal{E}$. The real power flow on line $\ell\in\mathcal{E}$ can be expressed in terms of the bus voltage magnitudes, angles, and the entries of the admittance matrix as
$p_{ij}$ $=|V_{i}||V_{j}|\Im(Y_{ij})\sin(\theta_{\ell})$
@@
where $M_{i}>0$ is the inertia coefficient of the synchronous generator connected at node $i\in\mathcal{V}_{G}$. The term $D_{i}>0$ represents the damping coefficient of the $i$th synchronous generator when $i\in\mathcal{V}_{G}$, the frequency-dependent component of the load connected to bus $i$ when $i\in\mathcal{V}_{L}$, and for droop-controlled inverter buses $i\in\mathcal{V}_{R}$, the total frequency dependence from the inverter’s droop and frequency-responsive loads (i.e., $R_{i}^{-1}+D_{i}^{\prime}$, where $R_{i}^{-1}$ is the droop coefficient and $D_{i}^{\prime}$ is the frequency-responsive component of the load) [1, 16], and $\omega_{i}$ is the natural frequency of the $i$th oscillator (this is the net real power injection at bus $i$).
With the integration of more renewable generation and the retirement of synchronous generation, we assume that the inertia coefficients $M_{i}$ are small compare damping term $D_{i}$ – this is typical for low inertia grids. [17] showed that under these assumptions, in the time-scale for synchronization, the network dynamics can be approximated by the dynamics of non-uniform Kuramoto oscillators where the model’s approximation error is $\mathcal{O}(\epsilon)$ and goes to zero as $t\rightarrow\infty$, $\epsilon>0$ is the singular perturbation parameter which typically is the worst-case choice of the ratio $M_{i}/D_{i}$. The slower dynamics of the network, that approximates the synchronization dynamics, is given by the reduced first-order model
@@
III Clustering
III-A Linearized Synchronization Dynamics
Let $|V^{*}|$ and $\delta^{*}$ denote the voltage magnitudes and phase angles at a frequency-synchronized solution (i.e. a power flow solution), if we assume that reactive power balance in the network can be controlled via local compensation, meaning that the voltage magnitudes $|V_{i}^{*}|$ can be assumed to be constant (not necessarily $1.0$ p.u.), the linearized voltage angle dynamics around $\delta^{*}$ is given by the weighted Laplacian flow,